### v1

# Yao Cheng, Qier Cui, Yunshi Lan, Xiang Li

Abstract
Peer review, as a cornerstone of scientific research, ensures the integrity and quality of scholarly work by providing authors with objective feedback for refinement. However, in the traditional peer review process, authors often receive vague or insufficiently detailed feedback, which provides limited assistance and leads to a more time-consuming review cycle. If authors can identify some specific weaknesses in their paper, they can not only address the reviewer’s concerns but also improve their work. This raises the critical question of how to enhance authors’ comprehension of review comments. In this paper, we present SEAGraph, a novel framework developed to clarify review comments by uncovering the underlying intentions behind them. We construct two types of graphs for each paper: the semantic mind graph, which captures the author’s thought process, and the hierarchical background graph, which delineates the research domains related to the paper. A retrieval method is then designed to extract relevant content from both graphs, facilitating coherent explanations for the review comments. Extensive experiments show that SEAGraph excels in review comment understanding tasks, offering significant benefits to authors.
[Uncaptioned image]   SEAGraph: Unveiling the Whole Story of Paper Review Comments
Jianxiang Yuthanks:   Equal Contribution, Jiaqi Tan11footnotemark: 1, Zichen Ding, Jiapeng Zhu, Jiahao Li, Yao Cheng, Qier Cui, Yunshi Lan, Xiang Li East China Normal University, Shanghai, China sea.ecnu@gmail.com
1 Introduction
In recent years, the number of academic publications has grown exponentially, creating a vast “sea of papers” Bornmann and Mutz (2015); Lin et al. (2023). Traditionally, authors rely on the peer review process to receive feedback on their manuscripts Lee et al. (2013); Björk and Solomon (2013). However, the review cycle typically requires several months or even longer, which is time-consuming Horbach and Halffman (2018). Meanwhile, the large volume of submissions results in uncertain review qualities Bornmann and Mutz (2015), often resulting in ambiguous or overly brief comments that are challenging to explain Verma et al. (2022). For example, the statement that “the method is limited” is very vague without any details provided. In such cases, authors may be confused about which parts of the paper to be revised, leading to the difficulty in effectively addressing the reviewers’ issues or fully identifying the shortcomings of the paper. If they receive detailed suggestions, they may benefit more during this review process. Therefore, it is essential to understand the vague comments with more details provided, thereby helping the authors to improve their paper quality (as shown in Fig. 1).
Figure 1: SEAGraph  can help authors better understand reviewers’ comments by providing detailed insights and evidence.
Currently, Large Language Models (LLMs) have shown powerful text comprehension and generation capabilities Achiam et al. (2023); Wei et al. (2022), offering new directions for revealing the underlying intentions behind each review comment. A straightforward approach is to provide LLMs with both the comment and the corresponding paper. Yet, it is usually difficult to feed an entire paper into LLMs for identifying key points, as review comments typically focus on specific aspects rather than the entire paper. Another alternative approach is using RAG (Retrieval-Augmented Generation) Cheng et al. (2024); Jiang et al. (2023), which enhances reasoning by retrieving the most relevant passages from lengthy texts based on the query. Nevertheless, the information retrieved by RAG tends to be fragmented, lacking clear logic Cao et al. (2024). In contrast, review comments are given based on the coherent logical structure formed when reviewers read the paper, which cannot be easily captured by fragmented segments. Recently, the success of GraphRAG Edge et al. (2024), which splits lengthy texts into discrete chunks and hierarchically connects them, has inspired new directions. Similarly, papers are inherently structurally organized with sections and subsections provided. Therefore, we can format papers as structured graphs, from which logical chains can be extracted to facilitate a deeper understanding of review comments.
In this paper, we propose SEAGraph, a novel framework designed to uncover the intentions behind paper reviews and enhance the understanding of review comments. We construct two distinct graphs for each reviewed paper: a semantic mind graph and a hierarchical background graph. Building upon the principles of the mind map D’Antoni and Zipp (2006), which employs visuospatial orientation to integrate information, we introduce the semantic mind graph to facilitate deeper semantic connections and organization of key knowledge points. In addition, the hierarchical background graph connects various related papers based on the themes of the paper, thereby simulating its research context. After the construction of the two graphs, we design a tailored retrieval method to extract the most relevant content from both graphs in response to each review comment. The extracted content is subsequently fed into LLMs to generate coherent and logical arguments that explain the reviewer’s comments. Overall, our contributions are summarized as follows:
  • We introduce a novel framework SEAGraph, which pioneers the field of review comment understanding.
  • We construct a semantic mind graph and a hierarchical background graph for a paper, capturing its deep semantics and related domain knowledge.
  • We conduct extensive experiments to validate the effectiveness of our framework, which can help authors improve the quality of their papers.
Figure 2: The overall framework of SEAGraph consists of the construction of the semantic mind graph and the hierarchical background graph, along with the corresponding retrieval module. The final retrieved content is fed into LLMs for review comment understanding.
2 Related Work
2.1 Retrieval-Augmented Generation
RAG (Retrieval-Augmented Generation) improves the generation performance of LLMs by incorporating external knowledge Lewis et al. (2020). Initially, naive RAG approaches follow a process including indexing, retrieval, and generation Li et al. (2022). Advanced RAG frameworks focus on enhancing the retrieval quality during the pre-retrieval and post-retrieval phase like query rewriting, query expansion, and chunk reranking  Ma et al. (2023); Peng et al. (2024b); Zheng et al. (2023). Furthermore, some modular RAG approaches put forward new modules or pipelines to enhance the retrieval capability and alignment with task-specific requirements Yu et al. (2022); Shao et al. (2023). Despite these advancements, RAG faces challenges in handling query-focused summarization tasks when queries target entire text corpora Cao et al. (2024). GraphRAG emerges as an innovative solution to address this challenge Peng et al. (2024a). Edge et al. (2024) establish logical relationships between segments by connecting chunks or communities through a hierarchical structure. Wu et al. (2024a) and Sepasdar et al. (2024) construct specialized knowledge graphs, extending GraphRAG to the medical and soccer domains. In this work, we construct two logically connected graphs for the paper, leveraging the strengths of GraphRAG to better address the review comment understanding tasks.
2.2 Large Language Models in Peer Review
Recently, Large Language Models (LLMs) have made remarkable progress in text generation tasks Zhao et al. (2023); Ouyang et al. (2022); Luo et al. (2024), prompting researchers to explore new opportunities in the field of peer review  Li et al. (2024b); Checco et al. (2021). A significant focus has been placed on generating automated reviews to enhance the quality of academic papers Gao et al. (2024). For example, Liu and Shah (2023) and Liang et al. (2023) customize prompts to guide GPT-4 in generating scientific feedbacks. Yu et al. (2024a) and Wei et al. (2023) respectively employ supervised fine-tuning and continuous pretraining to align LLMs with human reviews. Expanding on this research, Jin et al. (2024) employ LLMs to simulate the entire review process, thereby revealing the impact of various factors on academic evaluation. Subsequently, Ye et al. (2024) highlight the risks associated with using LLMs for automated peer reviewing. Building on these insights, Yu et al. (2024b) delve into the challenge of distinguishing between peer reviews authored by humans and those generated by LLMs. However, both human reviews and generated reviews may suffer from issues such as ambiguity or brevity, causing confusion for authors. Our work aims to leverage LLMs to understand the intent of review comments, thereby assisting authors in polishing their papers.
3 SEAGraph
This section provides a detailed description of each module in SEAGraph, and the overall framework is in Figure 2.
3.1 Review Comments Understanding Task
Paper Processing.
Our dataset consists of the PDF versions of academic papers and their corresponding review comments. We begin by utilizing Nougat Blecher et al. (2023) as the parser, a model built on the Visual Transformer architecture specifically tailored for extracting information from academic documents. Then, we construct the Semantic mind graph and the Hierarchical background graph for each reviewed paper, denoted as $\mathcal{G}_{S}(\mathcal{V}_{S},\mathcal{E}_{S})$ and $\mathcal{G}_{H}(\mathcal{V}_{H},\mathcal{E}_{H})$, where $\mathcal{V}$ and $\mathcal{E}$ represent nodes and edges, respectively.
Review Comments Extraction.
We input the entire review into LLM to extract individual comments under three textual sections: “Strengths, “Weaknesses”, and “Questions.” Then, all review comments are defined as a query set $\mathcal{Q}$, where $q\in\mathcal{Q}$ represents a single comment.
Our Goal.
Our goal is to retrieve subgraphs from two graphs $\mathcal{G}_{S}$ and $\mathcal{G}_{H}$, which represent evidence relevant to a review comment $q$. These pieces of evidence, along with the comment, are then provided to LLMs to help understand the details of the review. By generating a logical chain, SEAGraph enables authors to better understand reviewers’ perspectives and proceed with subsequent research more effectively.
3.2 Semantic Mind Graph Construction
A paper is structurally organized into different sections, while key points of the paper may be scattered across various parts. The entire paper can be structured like a mind graph, where content progressively branches out from different paragraphs. Our goal is to construct a semantic mind graph to model the writing logic of a paper. We next detail the main steps.
Paper Chunking.
We first use the Spacy library AI (2017) to break a full paper down to sentence level, allowing us to assess the relevance between sentences and decide whether adjacent sentences should be merged into chunks. Specifically, we utilize Sentence-BERT Reimers and Gurevych (2019) to encode both sentences and chunks, and design a semantic relevance measure to determine whether the current chunk is related to the next sentence. Subsequently, we place the first sentence $s_{1}$ into the initial chunk $c_{\text{current}}$. For each subsequent sentence $s_{i}$, we compute the embedding similarity between $s_{i}$ and $c_{\text{current}}$. If the similarity exceeds a threshold $\theta_{1}$, we merge the sentence into the current chunk by:
$c_{\text{current}}\leftarrow c_{\text{current}}\cup s_{i}\quad\text{if }\mathit{f}(h_{s_{i}},h_{c_{\text{current}}})\geq\theta_{1},$
where $h_{s_{i}}$ and $h_{c_{\text{current}}}$ represent the embeddings of the $i$-th sentence and the current chunk, respectively. Here, $\mathit{f}(\cdot,\cdot)$ is the similarity function (e.g. cosine similarity). If the similarity is less than the threshold, we end the current chunk and start a new one: $c_{\text{new}}\leftarrow s_{i}$. Then we repeat the above steps to divide a paper into chunks. Additionally, a maximum chunk size is also set to prevent excessive imbalance in the length of different chunks. Finally, the chunk nodes of the paper can be represented as $\mathcal{V}_{S}=\{c_{a},c_{1},c_{2},\dots c_{n}\}$, where $c_{a}$ denotes the abstract node of the paper. In this way, the content within the same chunk is closely related.
Chunk Linking.
Given the segmented chunks, the next step is to link them based on their contextual and semantic relationships. Generally, chunks within the same (sub)section often share the same topic. For example, the “Method” section describes the paper’s research methodology, while the “Experiments” section validates the proposed method through experimental design. Therefore, adjacent chunks in a sequential order are probably highly correlated and we call this contextual correlation. We can establish connections between them by setting $e_{c_{a},c_{1}}=1$ and $e_{c_{i},c_{i+1}}=1,\forall i\in[1,n-1]$ where $e$ denotes the edge between two chunk nodes. This approach also helps mitigate issues that sentences with high semantic relevance may be split into different chunks due to chunk size constraints.
Further, the authors may not simply follow a linear mind in organizing the paper. As shown in Figure 3, the “Introduction” section of a paper often lays the groundwork for understanding the problem, which is further elaborated in the “Method” section with detailed descriptions of the proposed approach or framework. Subsequently, the “Experiments” section validates these methods through practical evaluations. If these segments are extracted individually, they can still form a coherent logic. We call this semantic correlation. To capture their correlations, we compute the semantic similarities between different chunks and set a threshold $\theta_{2}$ to connect highly similar chunks:
$e_{c_{i},c_{j}}=1\quad\text{if }\mathit{f}(h_{c_{i}},h_{c_{j}})\geq\theta_{2}$ (1)
Finally, the paper is transformed into a semantic mind graph, where linked chunks represent either contextual proximity or semantic similarity.
Figure 3: Chunks from different sections can form a coherent logic chain.
3.3 Hierarchical Background Graph Construction
To effectively review a paper, a reviewer not only needs a deep understanding of the content but also a solid grasp of the knowledge in corresponding fields. Therefore, we construct a background graph with hierarchical relationships to simulate reviewers’ domain knowledge. This graph is organized into a three-layer structure: the themes of the reviewed paper, the abstracts of relevant papers, and the semantic mind graph for each individual paper.
Cited Paper Search.
We first locate the cited papers in the “Related Work” section of the reviewed paper, which represent the existing research achievements in the field. Subsequently, we extract the publication details of these papers from the “Reference” section.
Theme Summarization.
We next crawl the PDFs of referenced papers from Google Scholar, parse them into markdown format, and extract the abstracts and titles of each paper. Then, we feed them into LLM to summarize multiple themes and assign corresponding papers to each theme. In this way, we obtain a theme set related to the reviewed paper, denoted as $\mathcal{T}=\{t_{1},t_{2},\dots,t_{m}\}$ and $t$ refers to the descriptive summarization of a theme node.
Complementary Papers.
The authors may not always reference all foundational or cutting-edge papers in the field. Therefore, we aim to enrich the paper by incorporating both fundamental and recent studies within the research domain. Based on the extracted themes, we search and crawl the most popular and recent papers related to these themes from Google Scholar to enrich the background graph in terms of breadth and timeliness. After identifying the relevant papers related to the paper, we apply the method from Section 3.2 to construct a semantic mind graph for each of them.
Hierarchical Linking.
Now, for a reviewed paper, we construct its hierarchical background knowledge graph based on theme nodes, abstract nodes, and semantic mind graphs. The first level includes multiple theme nodes, each corresponding to a thematic description that encapsulates the research topics. The second level connects these theme nodes to abstract nodes, where each abstract serves as a concise summary of a paper, representing its key ideas and maintaining a direct association with its respective theme. The third level extends from the abstract nodes to semantic mind graphs, which provide fine-grained information, offering a deeper insight into the paper’s structure and details. This hierarchical design clearly delineates the logical progression from themes to papers and further to detailed information, forming a systematic framework for representing the research background.
3.4 Semantic Mind Graph Retrieval
We next introduce retrieving the semantic mind graph based on review comments and obtaining the relevant supporting texts. Given a review comment as a query, we first calculate the probability distribution of the query over the semantic mind graph by calculating the textual similarity between the query and each chunk node $c_{j}$:
$P(c_{j})=\frac{\mathit{f}(h_{q},h_{c_{j}})}{\sum_{i=1}^{n}\mathit{f}(h_{q},h_{c_{i}})},$ (2)
where $P(c_{j})$ represents the probability distribution over the nodes in the semantic mind graph, $\mathit{f(\cdot,\cdot)}$ is the similarity function and $h_{q}$ denotes the embedding of the query.
Then we iteratively retrieve chunk nodes that can help explain the review comments. We start by randomly sampling $k$ chunk nodes based on the probability computed in Eq. 2 and add them to an empty node set $\bar{\mathcal{V}}$. After that, we explore the one-hop neighbors of these newly sampled chunk nodes and add them to $\bar{\mathcal{V}}$. Given a chunk node $c_{i}\in\bar{\mathcal{V}}$, suppose chunk node $c_{j}$ as its one-hop neighbor and we calculate:
$\text{score}_{i}(h_{c_{j}})=\alpha\cdot P(c_{j})+\mathit{f}(h_{c_{i}},h_{c_{j}}),$ (3)
where $\mathit{f}(h_{c_{i}},h_{c_{j}})$ denotes the cosine similarity between the embeddings of the two chunks and $\alpha$ is a hyperparameter to control term balance. In Eq. 3, the score is used to measure the relevance between query and chunk $c_{j}$. In particular, the second term calculates the similarity between chunk nodes $c_{i}$ and $c_{j}$, which indirectly reflects the relation between query and chunk $c_{j}$ via chunk $c_{i}$. After the scores are computed, we select chunk nodes with the highest scores and further add them into $\bar{\mathcal{V}}$. We repeat the above process to retrieve more chunk nodes. Finally, all the chunk nodes in $\bar{\mathcal{V}}$ constitutes a subgraph that is relevant to the given review comment.
Method Relevance Clarity Criticality Novelty Practicality Persuasiveness Avg.Rank
DirectInfer 3.09 2.94 2.97 3.53 3.16 2.63 3.05
RAG-naive 2.72 2.78 3.22 2.91 2.72 3.16 2.92
RAG-SMG 2.63 2.97 2.50 2.41 2.91 2.78 2.70
RAG-HBG 4.41 4.13 4.25 4.22 4.16 4.31 4.25
SEAGraph 2.16 2.19 2.06 1.94 2.06 2.13 2.09
Table 1: Human evaluation results for Review Comments Understanding (Rank-Based $\downarrow$). We highlight the best score on each dataset in bold and the runner-up score with an underline.
3.5 Hierarchical Background Graph Retrieval
To further mine the background knowledge related to a paper, we conduct an in-depth hierarchical background graph retrieval based on the review comment and the corresponding semantic mind subgraph obtained in the previous section. The hierarchical retrieval process refines from (1) theme level to (2) abstract level, and finally to (3) chunk level, ensuring background knowledge obtained at different levels of granularity. We apply scoring method in Eq. 3 to the three levels of nodes and select the nodes with higher scores.
We first begin the retrieval process at the theme level, aiming to extract themes related to a review comment. For example, if a reviewer questions whether the proposed method shares similarities with certain techniques in the fields of computer vision (e.g., contrastive learning in CV), we retrieve the corresponding descriptions of theme nodes to align with the reviewer’s concerns from a broader perspective.
Based on the theme-level information, we then proceed to the abstract-level retrieval, focusing on papers related to the identified themes. Note that abstracts summarize key research questions, methodologies, and conclusions, providing a concise yet comprehensive overview. Therefore, abstract-level retrieval is particularly useful for understanding review comments that compare the originality of the proposed approach with existing methods.
Finally, we step into the chunk-level retrieval. Chunk nodes include detailed information, such as experimental setups and results. They can be used to better understand review comments, thereby providing details to revise the paper.
After retrieval across all three granularity levels, the top-$k$ relevant nodes are ranked and selected to ensure that the most pertinent evidence is used for understanding the review comments. More implementation details are provided in the Appendix B.
Upon retrieving informative nodes from both graphs, we feed the corresponding text along with the query into LLM to generate an explanation for the review comment.
4 Experiments
Figure 4: Automated evaluation results for Review Comments Understanding, measured by Win-Rate ($\uparrow$).
4.1 Experimental Setup
Datasets.
We select 20 papers and their corresponding reviews from ICLR over the past five years through OpenReview [^1] https://openreview.net/. Each paper was reviewed by 3 to 8 reviewers, with a total of 82 reviewers. We divide the review comments and exclude those longer than 200 characters, which are considered relatively complete. In the end, 284 shorter comments were retained. These papers cover diverse fields such as Artificial Intelligence (AI), Machine Learning (ML), Natural Language Processing (NLP), and Computer Vision (CV). The selected papers exhibit a wide range of citation counts, including both highly cited and less-cited works, as well as recent studies and earlier publications.
Baseline Methods.
To validate the effectiveness of SEAGraph in terms of graph construction and retrieval, we compare it with the following two categories of baseline methods: (1) Direct inference methods: DirectInfer takes the review comment and the full parsed paper as input to directly reason about the understanding of each review comment. (2) RAG-based methods: RAG-naive computes the similarity between each review comment and every chunk of the paper, selecting the top-$k$ chunks to combine with the review comment as input; RAG-SMG utilizes only the construction and retrieval of the Semantic Mind Graph; RAG-HBG relies solely the construction and retrieval of the Hierarchical Background Graph.
We conduct experiments using the same open-source base model Ministral-8B-Instruct-2410222https://mistral.ai/ and the same prompt (Table 3) as input to LLMs.
Evaluation Protocol.
Since the task of reivew comments understanding lacks a definitive ground truth and exhibits significant diversity in generated content, we design two evaluation methods: human evaluation and automated evaluation. For human evaluation, we engage 40 experts from various fields, each responsible for one paper to assess the quality of the generated understanding of the review comments. For automated evaluation, given the powerful text comprehension capabilities of LLMs to play as a judge Li et al. (2024a), we employ gpt-4o-2024-11-20333https://openai.com/ as the evaluation model to provide objective judgments for the task.
4.2 Main experiments
Evaluation Metric.
There are four main assessment metrics for human and automated evaluation: (1) Relevance: Assesses the alignment between the provided evidence and the review comments. (2) Clarity: Evaluates how clearly and effectively the information is presented for ease of understanding. (3) Criticality: Examines the depth of analysis and the extent to which the feedback reflects constructive thought. (4) Novelty: Measures the inclusion of fresh insights or new evidence. Additionally, two customized human-centric metrics are tailored for evaluation: (1) Persuasiveness: Focuses on the logical reasoning and the ability of arguments to persuade. (2) Practicality: Gauges the usefulness and applicability of the information for authors. The specific meaning of six metrics is provided in Appendix A.
Figure 5: Human evaluation results for Key concerns in Reviews. ($\uparrow$).
Human Evaluation Results.
In the human evaluation process, experts are invited to rank the review understanding results produced by five different methods based on six aforementioned metrics. The results are summarized in Table 1, where the scores indicate the average ranking over all samples. We can observe that: (1) SEAGraph achieves the highest performance across all metrics, underscoring the effectiveness of our framework. (2) In terms of average rankings, RAG-SMG performs better than RAG-naive, which in turn outperforms DirectInfer. These findings strongly support our motivation that logically retrieved content significantly improves the ability of LLMs to understand review comments. (3) Although RAG-HBG performs poorly—mainly because its retrieved content consists only of background knowledge without incorporating the internal knowledge of the paper—SEAGraph enhances the LLM’s ability to understand the paper by integrating external background knowledge on top of RAG-SMG, enabling it to extract more meaningful information for reasoning. Consequently, constructing the semantic mind graph and the hierarchical background graph can provide valuable support for understanding paper reviews from different perspectives.
Automated Evaluation Results.
In the automated evaluation, due to the excessive total text length generated by the five methods, we employ a pairwise ranking approach for assessment. Figure 4 presents the results for review comments with a length of less than 100, while the results and analysis for lengths between 100 and 200 are provided in Appendix C. The values in the heatmap represent the win rate of the method shown on the vertical axis over the method on the horizontal axis. From the figure, it can be observed that: (1) SEAGraph consistently performs as the optimal method across the four metrics. (2) In terms of Criticality and Novelty metrics, SEAGraph significantly outperforms other methods, indicating its ability to provide more innovative evidence and conduct deeper analysis for review comments. (3) In most cases, RAG-SMG consistently ranks second. Although it performs slightly worse than RAG-naive in terms of relevance, it surpasses RAG-naive across all other metrics. This suggests that while RAG-SMG captures slightly less relevant aspects of evidence, its ability to make logical connections greatly enhances the reasoning power of LLM. These findings highlight the crucial role of modeling academic papers as semantic mind graphs to capture the paper’s underlying structure and logic. (4) Notably, for novelty, RAG-HBG performs better than all methods except SEAGraph, as it introduces multiple perspectives beyond the reviewed paper.
Overall, both human and automated evaluations yield consistent conclusions, providing strong evidence that SEAGraph is capable of generating valuable insights for interpreting review comments.
4.3 Quantitative Evaluation of Retrieval
In this section, we automate the evaluation of retrieved content based on five distinct metrics: (1) Relevance, (2) Specificity, (3) Novelty, (4) Logic, (5) Explainability. The detailed meanings are shown in Appendix A. Due to the extensive length of the retrieved content, both the content and corresponding review comment are directly input into gpt-4o for scoring. The results, as presented in the Table 2, demonstrate that RAG-SMG outperforms other methods in the majority of cases. Specifically, while RAG-SMG scores slightly lower than RAG-naive on the Relevance metric, this discrepancy may be attributed to the retrieval of chunk nodes with marginally lower semantic similarity through the logical structure of the semantic mind graph. Conversely, the external background knowledge retrieved by RAG-HBG exhibits limited relevance to the review comment, leading to consistently lower scores across various evaluation metrics. In summary, the information retrieved via the semantic mind graph demonstrates greater contextual significance and utility.
4.4 Key Concerns in Reviews
In this section, we consider the issues highlighted in the review comments as the key concerns of the paper. We first use LLMs to summarize all the reviews of each paper, identifying the concerns most frequently mentioned by the reviewers. Then, we employ six metrics consistent with those in Section 4.2 and perform a human evaluation to quantify and compare the effectiveness of  SEAGraph in understanding and explaining these concerns with other methods. The results are shown in Figure 5. It can be seen that: SEAGraph consistently achieves best scores across all evaluation metrics, particularly excelling in persuasiveness and practicality when addressing the key concerns identified in the reviews, indicating a greater alignment with human preference. RAG-SMG also shows strong performance in all metrics, as the concerns of reviewers are closely aligned with the content of the paper. In contrast, RAG-HBG has a poor performance since the supplemental background knowledge is less helpful to the key concerns. In conclusion, both SEAGraph and RAG-SMG show superior performance in this experiment, garnering higher recognition for their practicality and persuasiveness from human evaluators, further confirming the superiority of our framework in assisting authors to understand reviews.
Metric $\uparrow$ RAG-naive RAG-SMG RAG-HBG
Relevance 6.46 6.45 4.75
Specificity 5.74 5.84 4.23
Novelty 5.06 5.10 3.68
Logic 6.57 6.63 4.82
Explainability 5.57 5.62 3.94
Table 2: Automated evaluation results for retrieval.
5 Conclusion
In this paper, we present SEAGraph, a novel framework designed to bridge the gap between reviewers’ comments and authors’ understanding. By constructing two distinct graphs—the semantic mind graph, which captures the authors’ thought process, and the hierarchical background graph, which encapsulates the research background—the framework effectively models the context of a reviewed paper. The well-designed retrieval method ensures that relevant content from both graphs is used to generate coherent and logical explanations for review comments. SEAGraph not only enhances the clarity of reviews but also empowers authors to understand reviewer concerns more effectively, improving the quality of academic publications.
In a nutshell, we sincerely hope that our work does empower authors to not only gain a deeper understanding of reviews feedback but also elevate the quality of their papers, ultimately expediting both the advancement of research and the efficiency of the submission process.
Limitations
SEAGraph is designed to assist authors in comprehending review comments during the peer review process, with a particular emphasis on the review-comment understanding stage. Here we elaborate on some of these constraints, along with intriguing future explorations.
Rebuttal mechanism.
The rebuttal mechanism, where authors respond to reviewers’ concerns and engage in further discussion, also plays a critical role in improving the paper Jin et al. (2024). The success of multi-agent systems in executing complex tasks presents a promising opportunity Wu et al. (2023, 2024b). In future research, we will explore simulating the rebuttal process through multi-agent communication, aiming to further bridge the understanding gap between reviewers and authors in papers and comments, thus advancing the rebuttal mechanism.
Enhancing Pipeline Stability.
As an integrated pipeline, SEAGraph involves a relatively complex process. Certain components, such as the processes of searching for and downloading relevant papers, are dependent on network conditions. To address this, we aim to continuously refine and optimize the underlying code, ensuring the robustness and stability of these technical operations while improving their overall efficiency.
Privacy Challenges in Benchmarking.
We plan to establish a standardized benchmark for the field of peer review to help standardize and advance research practices in this area. Currently, platforms such as OpenReview provide a wealth of publicly available papers and review data, offering valuable resources for studying the mechanisms and effectiveness of peer review. However, there are certain limitations to the use of these data, primarily concerning privacy protection. While the identities of authors are ostensibly anonymized, the peer review process allows senior roles such as Program Chairs and Area Chairs to access the actual identities of both authors and reviewers. This potential breach of privacy may pose challenges to the objectivity, fairness, and ethical considerations of related research.
Evaluation Metrics.
Although human evaluation and GPT-based evaluation can reflect the strengths and weaknesses of a model to some extent, they often involve significant subjectivity and lack consistency. This issue becomes particularly pronounced in open-ended generation tasks, where differences in standards and preferences among evaluators may lead to inconsistent results. Therefore, establishing a comprehensive, unified, and reproducible quantitative standard is crucial for more objective and fair assessment of model performance. Such a standard not only helps to minimize the influence of human bias but also provides more actionable feedback for subsequent model optimization and improvement.
Ethics Statement
This work seeks to assist authors in better comprehending review comments. We do not intend to suggest that some reviews are inherently of low quality or unhelpful. Instead, we appreciate that clearer and more comprehensible review comments can more effectively fulfill the primary objective of the peer review process—namely, to offer objective evaluations and constructive feedback aimed at improving the manuscript. Recently, some academic conferences have introduced AI-assisted review bots to standardize reviewers’ feedback. Through this work, we aim to benefit authors by enhancing their understanding of review comments, while also encouraging reviewers to consider the clarity of their feedback and strive for higher-quality reviews. Ultimately, we seek to foster a healthier and more harmonious academic interaction environment.
Appendix A More details of SEAGraph
Figure 6: The example of SEAGraph.
Prompt.
In Table 3, we present the instruction designed to generate content for understanding review comments that conform to the specified format based on the retrieved content. We require LLMs to output several evidence before the summary.
The specific meaning of metrics.
The metrics for evaluating the generation of understanding review comments and the retrieval content are shown in Table 4 and Table 5, respectively.
SEAGraph results
Figure 6 shows an example of the explanation for a review comment generated by SEAGraph. For privacy concerns, both the review comment and the generated content have been processed. As shown in the figure, the review comment points out that the paper only conducts experiments in its own designed experimental settings, and suggests that comparing the proposed method with other libraries would help demonstrate its validity. SEAGraph locates, through the retrieval of the semantic mind graph, that the paper mentions only part of the workloads in the “Experiments” and “Related Work” sections, highlighting that, although the method is effective, it does not compare with some of the libraries or algorithms mentioned. It also points out in the “Conclusion” that the method’s validity could be verified by comparing it with more real-world applications. Additionally, SEAGraph, through the retrieval of hierarchical background knowledge, mentions other papers that have conducted such comparisons in their experiments. Finally, in the summary, SEAGraph effectively consolidates the logic of the entire review comment, highlighting the missing experiments in the paper and referring to other papers’ experimental settings. This example demonstrates SEAGraph’s ability to generate explanations for review comments by constructing two graphs and retrieving relevant chunks.
Appendix B Hierarchical Background Graph Retrieval
In the Hierarchical Background Graph, to make background knowledge more aligned with the reviewed paper and the review comments, we conduct a retrieval process for external knowledge based on the review comments and semantic mind subgraph identified in Section 3.4. First, we compute the representation of the retrieved semantic mind subgraph $\bar{\mathcal{V}}$ using a pooling operation to obtain the subgraph representation. First, we obtain the subgraph representation by applying a pooling operation to the node representations of the retrieved semantic mind subgraph.
$h_{\bar{\mathcal{V}}}=\frac{1}{|{\bar{\mathcal{V}}|}}\sum_{c_{i}\in\bar{\mathcal{V}}}h_{c_{i}}$ (4)
On this basis, we design a three-layer retrieval framework, including theme-level, abstract-level, and chunk-level retrieval, to capture information at varying levels of granularity. The specific retrieval formulas are as follows:
(1) Theme level.
$\text{score}(h_{t_{i}})=\alpha_{t}\cdot\mathit{f}(h_{t_{i}},h_{q})+\beta_{t}\cdot\mathit{f}(h_{t_{i}},h_{S_{s}})$
(2) Abstract level.
$\text{score}(h_{a_{i}})=\alpha_{a}\cdot\mathit{f}(h_{a_{i}},h_{q})$ $+\beta_{a}\cdot\mathit{f}(h_{a_{i}},h_{S_{s}})$
$+\gamma_{a}\cdot\mathit{f}(h_{a_{i}},h_{t_{a}})$
(3) Chunk level.
$\text{score}(h_{c_{i}})$ $=\alpha_{c}\cdot\mathit{f}(h_{c_{i}},h_{q})+\beta_{c}\cdot\mathit{f}(h_{c_{i}},h_{S_{s}})$
$+\gamma_{c}\cdot\mathit{f}(h_{c_{i}},h_{t_{a}})+\delta_{c}\cdot\mathit{f}(h_{c_{i}},h_{a_{c})}$
Here, $\alpha$, $\beta$, $\gamma$, and $\delta$ represent the hyperparameters, while $t_{i}$, $a_{i}$, and $c_{i}$ correspond to the text of the theme description, the abstract, and the chunk, respectively. Additionally, $h_{q}$ denotes the embedding of the query.
Appendix C Supplement to Automated Evaluation
In Figure 7, we show a pairwise comparison of the text generated by the five methods in the automated evaluation. Compared to Figure 4 in the main text, here we show the results for samples of all lengths. The values in the heatmap represent the win rate of the method shown on the vertical axis over the method on the horizontal axis. From the figure, we can see that SEAGraph outperforms in most cases, only slightly falling short on the Relevance metric when compared to Infer. On the other hand, RAG-SMG lags slightly behind RAG-naive in terms of relevance, likely because longer review comments are already sufficiently detailed, leading to a minor disadvantage in argument relevance. However, both SEAGraph and RAG-SMG demonstrate superior performance on other metrics, proving that the evidence they provide are more effective and better support reasoning.
Prompt of the task of Review Comment Understanding. You are an experienced researcher with strong logical thinking and excellent reasoning skills. You will receive a paper along with a corresponding review comment. We have provided the key sections of the paper and the critical content from related work. The review comment is found between <REVIEW> and </REVIEW>, the key sections of the paper are between <PAPER HIGHLIGHTS> and </PAPER HIGHLIGHTS>, and the related work is found between <RELATED WORK> and </RELATED WORK>. Your task is to systematically find supporting evidence and construct a complete logical chain, thereby building a full reasoning chain to clarify why the reviewer made this comment. Please structure the reasoning chain as follows: - **Evidence 1 (specific section) **: <corresponding reasoning> - **Evidence 2 (specific section)**: <corresponding reasoning> - (more evidence if available) - **Summary**: <Logical reasoning based on evidence explaining the basis for the review comment.>
Table 3: Instruction for generating understanding content for review comments.
Metric Description
Relevance Is the evidence in the supplementary information highly relevant and closely aligned with the reviewers’ comments?
Clarity Is the supplementary information clearly articulated and easy to understand? Does it effectively explain the reviewers’ viewpoints and the supporting arguments?
Criticality Does the supplementary information provide an in-depth analysis and reflection on the reviewers’ feedback? Does it identify any limitations in the feedback and offer reasonable suggestions for improvement?
Novelty Does the supplementary information present unique insights or new evidence not mentioned in the original review, thereby enriching the depth and breadth of the content?
Persuasiveness Does the summary present the evidence in a compelling manner, demonstrating logical reasoning and effectively persuading the reviews of the core ideas with clarity and coherence?
Practicality Does the supplementary information provide direct assistance to the author?
Table 4: Metrics for Evaluating Supplementary Information.
Metric Description
Relevance How closely does the retrieved information relate to the topic of the paper review or the content of the paper?
Specificity Is the retrieved information detailed and specific? Can it effectively supplement the review content or provide new insights?
Novelty Does the retrieved information offer a new perspective or provide supportive evidence not mentioned in the review?
Logic Is the retrieved information consistent with the review content and the overall logic of the paper?
Explainability Can it effectively address the issues mentioned in the review or provide theoretical foundations or case studies to back up the review’s arguments?
Table 5: Metrics for Evaluating Retrieved Information.
Figure 7: Automated evaluation results for Review Comments Understanding, measured by Win-Rate ($\uparrow$).